OpenStudy (anonymous):

Please answer 12+ sqt y-2=4

OpenStudy (anonymous):

\[12+\sqrt{?}y-2=4\]

OpenStudy (anonymous):

Please someone help!!

OpenStudy (anonymous):

\[12+\sqrt{?}y-2=4\]